Circular No. 8411

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004du, 2004ek, 2004em, AND 2004en A. V. Filippenko, M. Ganeshalingam, and B. J. Swift, University of

California, Berkeley, write: “Inspection of CCD spectra (range 330–1000 nm), obtained on Sept. 24 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2004em (IAUC 8406) is of type II, shortly after the explosion. The object has a blue, nearly featureless continuum and clear hydrogen Balmer lines with P-Cyg profiles, but the emission component of Hα is considerably stronger than the absorption component. Weak He I 587.6-nm absorption is visible, as well. SN 2004en (IAUC 8406) is also of type II, but somewhat older, perhaps a few weeks after the explosion. SN 2004du (IAUC 8387) appears to be a normal type-II-P supernova with typical P-Cyg profiles; the narrow Hα emission line responsible for the previous classification of type IIn (IAUC 8388) was based partly on a superposed H II region. We confirm the suggestion that SN 2004ek (IAUC 8405) is a young type-II supernova with a blue continuum (IAUC 8409), but now the flux density shortward of 400 nm drops substantially. There are weak hydrogen Balmer and Fe II absorption lines, although the weak Hα line is almost entirely in emission.”

SUPERNOVA 2004eq IN ESO 404-G12 Y. Chen provides a corrected position for SN 2004eq (cf. IAUC 8407):

α = 21^{h}57^{m}09^{s}.46, δ = −34^{o}34′51′′.3 (equinox 2000.0), which is 23′′.1 east and 4′′.3 north of the center of ESO 404-G12 = PGC 67701. He adds that a bright nearby foreground star has position end figures 09^{s}.10, 48′′.1.

4U 0115+63 J. Tueller, Goddard Space Flight Center (GSFC), on behalf of the In-

FOCuS team (a collaboration between NASA/GSFC, Nagoya University, JAXA/ISAS, and the University of Pennsylvania), reports that observations of the transient x-ray pulsar 4U 0115+63 (cf. IAUC 3161, 7116) were made with a balloon-borne hard-x-ray focusing telescope (launched from Ft. Sumner, NM) for 3200 s starting at Sept. 17.19 UT. The pulsar is continuing an outburst first detected by RXTE/ASM during Aug. 27–Sept. 2 (cf. http://xte.mit.edu/XTE.html). A preliminary 20–50-keV flux of (3.18 _{−0.324}^{+0.375}) × 10^{−9} erg cm^{−2} s^{−1}, consistent with the RXTE/ASM data, was derived from data obtained during Sept. 17.20–17.22. A summed-epoch analysis yields a preliminary pulse period of 3.616 ± 0.001 s.

2004 September 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green